OpenStudy (anonymous):

The volume of cube P is 10. If the length of an edge of cube Q is 4 times the length of an edge of cube P, what is the volume of cube Q?

OpenStudy (anonymous):

640

OpenStudy (baldymcgee6):

For cube P: V = s^3 10 = s^3 side length = cbrt(10)

OpenStudy (baldymcgee6):

side length of Q = 4 * cbrt(10)

OpenStudy (anonymous):

what is the cube root of 10? I don't know if i'm doing it right on my calculator.

OpenStudy (baldymcgee6):

cbrt(10) = 2.15

OpenStudy (anonymous):

givenp^3=10 it is also given q=4p therefore q^3=(4p)^3=64p^3=64*10=640

OpenStudy (anonymous):

Sorry, but I still don't get it..

OpenStudy (baldymcgee6):

do you still need help?

OpenStudy (anonymous):

pl explain what is your doubt